Lead, Pb, is used in solders and as radiation shield. It adopts the face centered cubic unit structure with each of the atom having a diameter of 0.35 nanometer. Find the theoretical density of lead if the molecular weight of this metal is 207.2 g/mol.


1
Expert's answer
2020-09-29T06:52:05-0400
"V_{atom} = {4\pi \over 3}d^3 = {4\pi \over 3}(0.35 nm)^3 = 0.18 nm^3 = 0.18*10^{-27} m^3"

"V_{total} = V_{atom}N_A = 0.18*10^{-27}*6.02*10^{23} = 1.0836*10^{-4} m^3"

"d = {m \over V_{total}} = {207.2 g \over 1.0836*10^{-4} m^3} = 191.21*10^4 g/m^3 = 1912.1 kg/m^3"